Solve the system of equations however you see fit. You may use substitution or elimination.
y=x+1
y=-2x(minus)2
3x(minus)5y=1
2x+5y=9

Answer by mananth(16946)

y=x+1
y=-2x-2
...
x+1 = -2x-2 =y
x+3x=-2-1
4x=-3
/4
x= -3/4
plug value of x in y=x+1
y=-3/4 +1
y =1/4
...
3x-5y=1
2x+5y =9
add both equations
3x+2x=10
5x=10
/5
x=2
...
plug value of x in 2x+5y=9
4+5y=9
5y=9-4
5y=5
/5
y =1
...
m.ananth@hotmail.ca